OpenStudy (anonymous):

assume f(x)=2x+8 and g(x)=5. What is the value of (f+g)(3)?

OpenStudy (bahrom7893):

f+g = (2x+8)+5 = 2x+13 (f+g)(3) = 2*(3)+13=6+13=19